# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ Abstract
[inst1]organization=Space based Computing Systems Research Center, Zhejiang Lab,city=Hangzhou, postcode=311100, state=Zhejiang, country=China
@@
[inst2]organization=Department of Computer Science, The University of Hong Kong,city=Hong Kong SAR, country=China
@@ 1 Introduction
Image super-resolution (SR) involves reconstructing high-resolution (HR) images from low-resolution (LR) images [1]. As a fundamental computer vision task, SR has far-reaching applications in various domains, including remote sensing [2], medical imaging [3], and security investigation [4]. However, remote sensing images (RSI) often suffer from limited spatial resolution due to sensor constraints, such as resolution, cost, orbit altitude, and transmission bandwidth. Despite these limitations, high-quality HR images are essential for real-world remote sensing applications. To address this challenge, remote sensing image super-resolution (RSISR) aims to enhance the detailed information of RSI by generating HR images from LR images, as shown in Fig. 1. RSISR has been extensively applied in land cover mapping [5], agricultural monitoring [6], disaster assessment [7], and urban monitoring [8].
Typically, RSISR approaches are categorized as single-frame or multi-frame based on input count. Single-frame methods utilize a single LR image to reconstruct a HR output, whereas multi-frame techniques exploit multiple LR observations from diverse perspectives or spectral bands to enhance reconstruction fidelity. Multi-frame SR benefits from complementary information across temporal, multi-view, or multi-sensor data, enabling improved recovery of fine details and high-frequency components. However, practical deployment is constrained by the limited availability of multi-temporal or multi-sensor imagery [9] and the computational complexity of accurate image registration. Satellite-based applications, for instance, face challenges in acquiring multi-temporal data due to long revisit intervals. Consequently, several studies [10, 11, 12, 13] have focused on single-frame RSISR to circumvent these limitations. Motivated by these considerations, this survey primarily examines single-frame RSISR methodologies.
@@
Figure 1: The benefit of remote sensing images super-resolution. The first row shows the low-resolution images, while the second row shows the results of 4x super-resolution using the SRCNN method [32, 33]. As can be seen, the super-resolution images contain richer details and texture information.
Despite the numerous papers on RSISR methods published annually, there is a scarcity of comprehensive survey papers that thoroughly summarize these algorithms. While several excellent survey papers on RSISR exist [9, 14, 34, 35, 36, 2, 37, 38, 25], they remains limitations and challenges. For instance, early review papers primarily focused on conventional RSISR algorithms [9, 14], which are insufficient to meet current needs due to the lack of investigation of deep learning methods. Additionally, some surveys only cover specific types of methods, such as deep learning-based methods [35, 36, 2, 38] and GAN-based methods [25]. Liu et al. [34] conducted an excellent survey on RSISR, but it was published in 2021 and does not cover algorithms published from 2022 to 2025.
@@
In this paper, we address these gaps by presenting a detailed, comprehensive, and systematic review of RSISR methods developed over the past two decades, covering more than 300 references. Meanwhile, the background, taxonomy, recent development characteristics, datasets and evaluation metrics in RSISR are also summarized in detail. In summary, the main contributions of this paper lie in the following aspects:
  • 1.
    A comprehensive review of RSISR methods. We conduct a thorough survey on the development of RSISR algorithms in the past two decades, including methods, datasets, and evaluation metrics. For a clearer discussion, various RSISR algorithms are classified into different categories, and their representative algorithms are introduced in detail. To the best of our knowledge, this is the most comprehensive review of RSISR methods to date.
  • 2.
    A systematic summary of the development characteristics of RSISR methods. Recent development features are summarized based on an in-depth survey of more than 300 RSISR papers. Further, development trends and challenges can be extrapolated.
  • 3.
    A detailed discussion of future prospects. Based on a comprehensive investigation and practical experience, future prospects of RSISR methods are discussed. These insights are expected to encourage further exploration of fundamental challenges and inspire broader engagement in this rapidly evolving field.
The rest of paper is organized as follows. Section 2 introduces the background and taxonomy of RSISR methods. Section 3 and 4 elaborate on the supervised and unsupervised RSISR methods respectively, followed the recent development characteristics of the RSISR field in Section 5. Then, Section 6 analyzes the summary of datasets and evaluation metrics. Next, Section 7 discusses the performance evaluation methods for RSISR. and Section 8 discusses future prospects. Finally, Section 9 concludes this paper.
@@ 2.1 Background of RSISR
RSISR has evolved over two decades, encompassing diverse technical branches such as multispectral image super-resolution (MSISR) [39], hyperspectral image super-resolution (HSISR) [40] and pan-sharpening [41]. RSISR has emerged as a critical research field due to the inherent resolution limitations of satellite and aerial imaging systems, which directly affect the performance of downstream applications such as land-use monitoring [5], disaster assessment [7], and urban monitoring [8].
@@ 2.2 Taxonomy of RSISR Algorithms
Figure 3: Taxonomy of the existing RSISR methods.
RSISR algorithms can be classified in different ways. Based on the latest trends in image processing, this paper conducts a comprehensive survey of RSISR algorithms from unsupervised and supervised perspectives. Fig. 3 shows a taxonomy of existing RSISR methods in this paper. Specifically, unsupervised methods and supervised methods can been grouped into traditional methods and deep learning based methods, respectively. It should be mentioned that most existing traditional RSISR methods are unsupervised, which mainly focus on the intrinsic structural characteristics and prior knowledge of the image. In addition, most deep learning-based RSISR methods are supervised, which mainly rely on large-scale labeled data to learn more complex mapping. Furthermore, supervised traditional methods can be divided into sparse coding-based, neighborhood embedding-based and regression-based methods, while unsupervised traditional methods can be divided into interpolation-based, Projection onto Convex Sets (POCS)-based, Iterative Back Projection (IBP)-based and probability-based methods. According to the type of adopted model, supervised deep learning methods can be divided into CNN-based, GAN-based, Transfrmer-based, Diffusion-based, Mammba-based and hybrid model-based methods. In contrast, unsupervised deep learning methods can be divided into self-supervised learning, contrastive learning, generative methods, and zero-shot learning methods. An overview of representative RSISR methods is given in Table 1.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3 Supervised RSISR Methods
Benefiting from machine learning methods, supervised learning-based methods are outperforming traditional methods. Generally, the supervised SR algorithms require training with HR-LR image pairs to learn the mapping relationship from LR images to HR images. This section surveys the supervised RSISR algorithms in detail from two aspects: traditional methods and deep learning-based methods.
@@ 3.1.1 Sparse coding-based methods
Sparse coding is a learning-based method that represents a signal via an over-complete dictionary with a small number of non-zero coefficients. To the best of our knowledge, Yang et al. [115, 116] proposed the first sparse coding-based SR method, which mainly consists of three algorithmic steps: dictionary learning, sparse coding and HR reconstruction, as shown in Fig. 4. Specifically, an over-complete dictionary is first learned from the training patches, then the test patches are represented by sparse coefficients, and finally the weighted coefficients are used to reconstruct the HR image.
@@
Some studies based on sparse representation have been conducted to reconstruct the HR images in remote sensing. For example, Zheng et al. [117] proposed an improved sparse representation method for joint RSISR and denoising. Specifically, they employed Batch Orthogonal Matching Pursuit (Batch-OMP) for LR encoding and achieving HR patch reconstruction via jointly-trained LR-HR dictionaries with sparse representation consistency. In addition, Zhang et al. [118] divides the over-complete dictionary into the primitive dictionary pair and the residual dictionary pair. The former is used to generate the initial HR image from the LR image, and the latter is learned to reconstruct the residual loss information. Similar work in Gou et al. [119] was performed to improve performance, which presented a non-local paired dictionary learning model involving an estimated dictionary and a residual dictionary. Considering that the sparse coefficients of the observed images of LR and HR images can hardly be kept consistent, Shao et al. [18] proposed a novel coupled sparse autoencoder (CSAE) to learn the mapping relationship between LR and HR images. Besides, Deka et al. [82] proposed an adaptive joint sparse representation-based method for RSISR. The main novelty is that its dictionary learning and sparse reconstruction only depend on the input LR images. A sparse structural manifold embedding (SSME) method proposed by Wang and Liu [50] incorporated the geometric properties of images into the neighbors calculation to recover the structure and edge detail information of HR images.
@@
Although sparse representation-based methods are effective in RSISR tasks, they still face many challenges, such as low computational efficiency, high dependence on dictionaries and low robustness. In addition, with the massive increase in remote sensing data, the real-time requirements of RSISR are becoming increasingly higher, which greatly limits the development of sparse representation methods.
3.1.2 Neighborhood Embedding-based
Figure 5: Two different RSISR framework. Top row: NE-based methods; Bottom row: regression-based methods.
@@ 3.1.2 Neighborhood Embedding-based
Neighborhood embedding (NE) based methods consider that small patches of LR and HR images form low-dimensional nonlinear manifolds with uniform local geometry [120]. That is, as long as the number of samples is sufficient, the corresponding image in the HR domain can be reconstructed by weighted averaging the neighbors in the LR domain using the same weights. Generally, NE-based RSISR method consists of the following three steps: manifold learning, LR neighborhood reconstruction, and super-resolution generation, as shown in Fig. 5 (top row). Specifically, manifold learning calculates the reconstruction error by neighborhood weighting for HR and LR manifolds from a specific training set and performs a minimization optimization. Given a patch $p_{i}$ from the neighborhood $N_{i}$, the reconstruction error $\varepsilon$ of HR or LR manifolds can be expressed as:
${\varepsilon}_{i}=||{p_{i}}-\sum\limits_{{p_{i}}\in{N_{i}}}{{{\bf{w}}_{i,j}}\cdot{p_{j}}}|{|_{2}}$ (2)
@@
To the best of our knowledge, Chang et al. [121] first proposed the NE-based single image super-resolution method, which employed local linear embedding (LLE) [120] and achieved a good super-resolution performance. Specifically, they perform local compatibility and smoothness constraints between patches in the target HR image via overlap, demonstrating the flexibility and effectiveness of the proposed method. In some studies, feature extraction strategies are utilized to improve performance. For example, Chan et al. [42] proposed a NE-based super-resolution method by combining edge detection and feature selection, named NeedFS, which performs a specific selection of training patches based on joint features and edge detection. In addition, Xu et al. [122] proposed a two-direction self-learning SR method based on a random oscillation+horizontal propagation+vertical propagation strategy. Bevilacqua et al. [123] proposed a low-complexity SR method based on nonnegative NE, which utilized a compact and efficient feature representation strategy to reduce the algorithmic complexity. A work by Timofte et al. [45] combines sparse dictionary learning and anchored neighborhood regression for fast super-resolution tasks. In particular, they used global collaborative coding to reduce the super-resolution map to a pre-computed projection matrix, which is extremely effective in improving the algorithm speed. Yang et al. [46] proposed a dual-geometric neighbor embedding (DGNE) method for SR task, which employed the multiview features and local spatial neighbors to find manifold embedding for image.
@@
NE-based RSISR methods take into account local neighborhood information and can effectively restore high-frequency information of the image, especially in details such as textures and edges. In addition, it retains the structural information of the image and can better avoid over-smoothing. Compared with deep learning methods, the NE-based methods have lower computational complexity and are suitable for resource-constrained environments, especially for large-scale remote sensing data processing. Although the NE-based methods are effective in RSISR, it has obvious disadvantages. For example, NE-based methods focus on the local neighborhoods, and its performance may degrade when processing global information in a larger range. In addition, the choice of neighborhood size has a great impact on the super-resolution results. Specifically, an inappropriate neighborhood size may lead to information loss or overfitting.
3.1.3 Regression-based

### v3

@@ Abstract
[inst1]organization=Research Center for Space Computing System, Zhejiang Lab,city=Hangzhou, postcode=311100, state=Zhejiang, country=China
@@
[inst2]organization=School of Computing and Data Science, The University of Hong Kong,city=Hong Kong SAR, country=China
@@ 1 Introduction
Image super-resolution (SR) involves reconstructing high-resolution (HR) images from low-resolution (LR) images [1]. As a fundamental computer vision task, SR has far-reaching applications in various domains, including remote sensing [2], medical imaging [3], and security investigation [4]. However, remote sensing images (RSIs) often suffer from limited spatial resolution due to sensor constraints, such as resolution, cost, orbit altitude, and transmission bandwidth. Despite these limitations, high-quality HR images are essential for real-world remote sensing applications. To address this challenge, remote sensing image super-resolution (RSISR) aims to enhance the detailed information of RSI by generating HR images from LR images, as shown in Fig. 1. RSISR has been extensively applied in land cover mapping [5], agricultural monitoring [6], disaster assessment [7], and urban monitoring [8].
Typically, RSISR approaches are categorized as single-frame or multi-frame based on input count. Single-frame methods utilize a single LR image to reconstruct an HR output, whereas multi-frame techniques exploit multiple LR observations from diverse perspectives or spectral bands to enhance reconstruction fidelity. Multi-frame SR benefits from complementary information across temporal, multi-view, or multi-sensor data, enabling improved recovery of fine details and high-frequency components. However, practical deployment is constrained by the limited availability of multi-temporal or multi-sensor imagery [9] and the computational complexity of accurate image registration. Satellite-based applications, for instance, face challenges in acquiring multi-temporal data due to long revisit intervals. Consequently, several studies [10, 11, 12, 13] have focused on single-frame RSISR to circumvent these limitations. Motivated by these considerations, this survey primarily examines single-frame RSISR methodologies.
@@
[…]
Despite the numerous papers on RSISR methods published annually, there is a scarcity of comprehensive survey papers that thoroughly summarize these algorithms. As shown in Table 1, we analyzed and compared existing RSISR reviews across key dimensions such as methodological scope, categorization strategy, and coverage of emerging deep learning architectures. Although these survey works [9, 14, 34, 35, 36, 2, 37, 38, 25, 39] are excellent, they still face significant challenges and limitations as follows:
  • Limited Scope: Early review papers [14, 9] primarily focused on traditional RSISR algorithms, which are insufficient to meet current needs due to the lack of investigation of deep learning methods. Additionally, some surveys only cover specific types of methods, such as deep learning-based methods [35, 2, 38] and GAN-based methods [25].
  • Outdated Coverage: Liu et al. [34] conducted an excellent survey on RSISR, but it was published in 2021 and does not cover algorithms published from 2022 to 2025.
  • Inconsistent Categorization: Existing surveys lack a unified and inclusive framework for classifying the diverse range of RSISR algorithms.
  • Limited Emerging Architectures Coverage: None comprehensively covers the latest and impactful deep learning paradigms like Transformers, Diffusion model, and Mamba.
@@
Figure 3: Taxonomy of the existing RSISR methods.
Table 1: Comparison of existing RSISR Surveys. Only peer-reviewed surveys are summarized in this table.
     Surveys/Reference     Year     RSI     Methodological Scope     Categorization Strategy     Deep Learning
                        Transformer     Diffusion Model     Mamba
     Yang et al. [14]     2015     ✓     Traditional     Mathematical Modeling Paradigm     ✗     ✗     ✗
     Fernandez et al. [9]     2017     ✓     Traditional     Task-driven Functional Taxonomy     ✗     ✗     ✗
     Liu et al. [34]     2021     ✓     Traditional + Deep learning     Methodological Chronology Perspective     ✗     ✗     ✗
     Wang et al. [35]     2022     ✓     Deep learning     Architectural Design-centric Taxonomy     ✓     ✗     ✗
     Karwowska and Wierzbicki [36]     2022     ✓     Traditional + Deep learning     Platform-aware Operational Framework     ✗     ✗     ✗
     Wang et al. [2]     2022     ✓     Deep learning     Model Development Workflow     ✗     ✗     ✗
     Lepcha et al. [37]     2023     ✗     Traditional + Deep learning     Algorithmic Hierarchy and Taxonomy     ✓     ✗     ✗
     Chen et al. [38]     2023     ✓     Deep learning     Hyperspectral-Specific Representation     ✓     ✗     ✗
     Wang et al. [25]     2023     ✓     Deep learning     Generative Adversarial Modeling Paradigm     ✓     ✗     ✗
     Mekhlafi and Liu [39]     2024     ✗     Traditional + Deep learning     Distillation-based Optimization Taxonomy     ✗     ✗     ✗
     Ours     2025     ✓     Traditional + Deep learning     Learning Supervision Paradigm Taxonomy     ✓     ✓     ✓
In this paper, we address these critical gaps by presenting a detailed, comprehensive, and systematic review of RSISR methods developed over the past two decades, covering more than 400 references. Furthermore, we provide detailed summaries of the background, taxonomy, recent development characteristics, datasets, and evaluation metrics fundamental to RSISR. To the best of our knowledge, this work is the most comprehensive, up-to-date, and well-structured review on RSISR to date, effectively filling the identified gaps in the existing literature. The main contributions of this paper are summarized as follows:
  • A novel & inclusive categorization: We introduce the learning supervised classification paradigm to systematically categorize existing RSISR algorithms, providing a more unified and comprehensive framework than previous taxonomies.
  • An unprecedented temporal coverage: We encompass the latest research progress up to 2025, including very recent publications missed by prior surveys.
  • A pioneering coverage of emerging architectures: This is the first RSISR survey to thoroughly integrate and discuss the impact of the three most significant emerging deep learning architectures: Transformers, Diffusion models, and Mamba.
  • A systematic summary of the development characteristics of RSISR methods: Recent development features are summarized based on an in-depth survey of more than 400 RSISR papers. Further, development trends and challenges can be extrapolated.
  • A detailed discussion of future prospects: Based on a comprehensive investigation and practical experience, future prospects of RSISR methods are discussed. These insights are expected to encourage further exploration of fundamental challenges and inspire broader engagement in this rapidly evolving field.
The rest of this paper is organized as follows. Section 2 introduces the background and taxonomy of RSISR methods. Section 3 and 4 elaborate on the supervised and unsupervised RSISR methods, respectively. Section 5 surveys RSISR methods from a modality-specific perspective, followed by a summary of datasets and evaluation metrics in Section 6. Section 7 discusses the performance evaluation methods for RSISR. Next, Section 8 summarizes the recent development characteristics of the RSISR and Section 9 discusses future prospects. Finally, Section 10 concludes this paper.
@@ 2.1 Background of RSISR
RSISR has evolved over two decades, encompassing diverse technical branches such as multispectral image super-resolution (MSISR) [40], hyperspectral image super-resolution (HSISR) [41], and pan-sharpening [42]. RSISR has emerged as a critical research field due to the inherent resolution limitations of satellite and aerial imaging systems, which directly affect the performance of downstream applications such as land-use monitoring [5], disaster assessment [7], and urban monitoring [8].
@@ 2.2 Taxonomy of RSISR Algorithms
RSISR algorithms can be classified in different ways, such as model architecture-based, input data modality-based, application-oriented, processing scope-based, loss function-based, etc. Given its fundamental impact on model design, training requirements, and real-world applicability, this review adopts the supervised vs. unsupervised dichotomy as the main axis of classification. Other taxonomies, including those based on architecture, data modality, application focus, processing scope, and loss strategies, are discussed within each primary category to provide a multi-faceted understanding of RSISR methods. Fig. 3 shows a taxonomy of existing RSISR methods in this paper. Specifically, unsupervised methods and supervised methods can be grouped into traditional methods and deep learning based methods, respectively. It should be mentioned that most existing traditional RSISR methods are unsupervised, which mainly focus on the intrinsic structural characteristics and prior knowledge of the image. In addition, most deep learning-based RSISR methods are supervised, which mainly rely on large-scale labeled data to learn a more complex mapping. Furthermore, supervised traditional methods can be divided into sparse coding-based, neighborhood embedding-based, and regression-based methods, while unsupervised traditional methods can be divided into interpolation-based, Projection onto Convex Sets (POCS)-based, Iterative Back Projection (IBP)-base,d and probability-based methods. According to the type of adopted model, supervised deep learning methods can be divided into CNN-based, GAN-based, Transformer-based, Diffusion-based, Mamba-based, and hybrid model-based methods. In contrast, unsupervised deep learning methods can be divided into self-supervised learning, contrastive learning, generative methods, and zero-shot learning methods. An overview of representative RSISR methods are given in Table 2.
@@
   DRSEN [61]     2019     CNN     Optical     Residual squeeze-excitation channel attention
@@
   RSMAP [64]     2019     Probability theory     SAR     Rayleigh-based sparse Bayesian deconvolution     ✓     ✓
@@
   CDGANs [68]     2020     GAN     Optical     Coupled-discriminator GAN with dual-path adversarial learning
@@
   CGAN [71]     2021     GAN     Optical     Cascade generative adversarial networks with scene-content constraints
   MIP [72]     2021     Generative method     Optical     Unsupervised migration prior generative super-resolution     ✓         ✓(Pytorch)
@@
   SSRN [75]     2021     Self-Supervised     Hyperspectral     Self-supervised spectral-spatial residual fusion     ✓
@@
   Fusformer [79]     2022     Transformer     Hyperspectral     Transformer-based spatial residual estimation             ✓(Pytorch)
@@
   MA-GAN [26]     2022     GAN     Optical     Multi-attention GAN with pyramid-convolutional mechanisms             ✓(Pytorch)
@@
   RAN [88]     2023     Contrastive learning     Optical     Contrastive region-aware graph super-resolution
@@
   DRSR [90]     2023     Self-Supervised     Hyperspectral     Self-supervised degradation-guided adaptive super-resolution     ✓         ✓(Pytorch)
@@
   Self-FuseNet [91]     2023     Zero-Shot learning     Optical     Forward self-fusion single-image super-resolution     ✓
@@
   SCDM [94]     2024     Diffusion model     Hyperspectral     Spectral-cascaded Diffusion Model with Image Condition Guidance             ✓(Pytorch)
@@
   SRDSRAN [100]     2024     GAN     Optical     Stratified dense sampling residual attention with chain training             ✓(Pytorch)
@@
   ESTNet [101]     2024     Transformer     Optical     Efficient swin transformer with channel-group attention             ✓(Pytorch)
@@
   ConvMambaSR [105]     2024     Mamba     Optical     State-space and CNN integrated global-local fusion
@@
   EDRLN [107]     2024     Contrastive learning     Optical     Contrastive degradation-aware efficient super-resolution             ✓(Pytorch)
@@
   IRR-DiffSR [111]     2025     Diffusion model     Optical     Image reconstruction representation-guided efficient diffusion             ✓(Pytorch)
@@
   DRGAN [113]     2025     GAN     Optical     Dynamic-convolution self-attention dense GAN
@@ 3 Supervised RSISR Methods
Recent advances in machine learning have enabled supervised learning-based methods to surpass traditional approaches [116]. Generally, the supervised SR algorithms require training with HR-LR image pairs to learn the mapping relationship from LR images to HR images. In this section, we provide a comprehensive survey of supervised RSISR algorithms, categorizing them into two primary paradigms: traditional methods and deep learning-based approaches.
@@ 3.1.1 Sparse Coding-based Methods
Sparse coding is a learning-based method that represents a signal via an overcomplete dictionary with a small number of non-zero coefficients. To the best of our knowledge, Yang et al. [117, 118] proposed the first sparse coding-based SR method, which mainly consists of three algorithmic steps: dictionary learning, sparse coding, and HR reconstruction, as shown in Fig. 4. Specifically, an over-complete dictionary is first learned from the training patches, then the test patches are represented by sparse coefficients, and finally, the weighted coefficients are used to reconstruct the HR image.
@@
Some studies based on sparse representation have been conducted to reconstruct the HR images in remote sensing. For example, Zheng et al. [119] proposed an improved sparse representation method for joint RSISR and denoising. Specifically, they employed Batch Orthogonal Matching Pursuit (Batch-OMP) for LR encoding and achieving HR patch reconstruction via jointly-trained LR-HR dictionaries with sparse representation consistency. In addition, Zhang et al. [120] divided the over-complete dictionary into the primitive dictionary pair and the residual dictionary pair. The former is used to generate the initial HR image from the LR image, and the latter is learned to reconstruct the residual loss information. Similar work in Gou et al. [121] was performed to improve performance, which presented a non-local paired dictionary learning model involving an estimated dictionary and a residual dictionary. Considering that the sparse coefficients of the observed images of LR and HR images can hardly be kept consistent, Shao et al. [18] proposed a novel coupled sparse autoencoder (CSAE) to learn the mapping relationship between LR and HR images. Besides, Deka et al. [83] proposed an adaptive joint sparse representation-based method for RSISR. The main novelty is that its dictionary learning and sparse reconstruction only depend on the input LR images. A sparse structural manifold embedding (SSME) method proposed by Wang and Liu [51] incorporated the geometric properties of images into the neighbors calculation to recover the structure and edge detail information of HR images.
@@
Although sparse representation-based methods are effective in RSISR tasks, they still face many challenges, such as low computational efficiency, high dependence on dictionaries, and low robustness. In addition, with the massive increase in remote sensing data, the real-time requirements of RSISR are becoming increasingly higher, which greatly limits the development of sparse representation methods.
3.1.2 Neighborhood Embedding-based Methods
Figure 5: NE-based RSISR method
@@ 3.1.2 Neighborhood Embedding-based Methods
Neighborhood embedding (NE) based methods consider that small patches of LR and HR images form low-dimensional nonlinear manifolds with uniform local geometry [122]. That is, as long as the number of samples is sufficient, the corresponding image in the HR domain can be reconstructed by weighted averaging the neighbors in the LR domain using the same weights. Generally, the NE-based RSISR method consists of the following three steps: manifold learning, LR neighborhood reconstruction, and super-resolution generation, as shown in Fig. 5. Specifically, manifold learning calculates the reconstruction error by neighborhood weighting for HR and LR manifolds from a specific training set and performs a minimization optimization. Given a patch $p_{i}$ from the neighborhood $N_{i}$, the reconstruction error $\varepsilon$ of HR or LR manifolds can be expressed as:
[…]
@@
To the best of our knowledge, Chang et al. [123] first proposed the NE-based single image super-resolution method, which employed local linear embedding (LLE) [122] and achieved a good super-resolution performance. Specifically, they perform local compatibility and smoothness constraints between patches in the target HR image via overlap, demonstrating the flexibility and effectiveness of the proposed method. In some studies, feature extraction strategies are utilized to improve performance. For example, Chan et al. [43] proposed a NE-based super-resolution method by combining edge detection and feature selection, named NeedFS, which performs a specific selection of training patches based on joint features and edge detection. In addition, Xu et al. [124] proposed a two-direction self-learning SR method based on a random oscillation+horizontal propagation+vertical propagation strategy. Bevilacqua et al. [125] proposed a low-complexity SR method based on nonnegative NE, which utilized a compact and efficient feature representation strategy to reduce the algorithmic complexity. A work by Timofte et al. [46] combines sparse dictionary learning and anchored neighborhood regression for fast super-resolution tasks. In particular, they used global collaborative coding to reduce the super-resolution map to a pre-computed projection matrix, which is extremely effective in improving the algorithm speed. Yang et al. [47] proposed a dual-geometric neighbor embedding (DGNE) method for the SR task, which employed the multiview features and local spatial neighbors to find a manifold embedding for the image.
@@
NE-based RSISR methods take into account local neighborhood information and can effectively restore high-frequency information of the image, especially in details such as textures and edges. In addition, it retains the structural information of the image and can better avoid over-smoothing. Compared with deep learning methods, the NE-based methods have lower computational complexity and are suitable for resource-constrained environments, especially for large-scale remote sensing data processing. Although the NE-based methods are effective in RSISR, it has obvious disadvantages. For example, NE-based methods focus on the local neighborhoods, and their performance may degrade when processing global information in a larger range. In addition, the choice of neighborhood size has a great impact on the super-resolution results. Specifically, an inappropriate neighborhood size may lead to information loss or overfitting.
3.1.3 Regression-based Methods